trains a andb are traveling in the same direction on parallel tracks. train a is traveling at 80 miles per hour and train b is traveling at 100 miles per hour. train a passes a station at 3:10 pm. If train B passes the same station at 3:22 pm at what time will train b catch up to a.
Answer by ankor@dixie-net.com(6693)

trains a and b are traveling in the same direction on parallel tracks. train a is traveling at 80 miles per hour and train b is traveling at 100 miles per hour. train a passes a station at 3:10 pm. If train B passes the same station at 3:22 pm at what time will train b catch up to a.
:
from the given information, we know that train B is 12 min (.2 hr) behind
train A, when train A passes the station
:
The distance between the trains at this time: .2 * 100 = 20 mi
:
Let t = time required for train B to catch train a
:
write a distance equation: Dist = speed * time
:
Train b travel dist = Train a travel dist + 20 mi
100t = 80t + 20
100t - 80t = 20
20t =
t = 1 hr
:
3:10 + 1:00 = 4:10 pm, B catches A
:
:
Check solution by finding the distances (train b travels 20 mi more than train b)
100 * 1 = 100 mi
80 * 1 = 80 mi
-------------------
difference = 20 mi